# Scientific Large Language Models: A Survey on Biological & Chemical Domains

### v1

@@ 7. Multi-modal Scientific Large Language Models
MoleculeSTM (Liu et al., 2023a) 2023.12 252M GraphMVP (Liu et al., 2021b), SciBERT Chemformer (Irwin et al., 2022) PubChemSTM
@@
ProteinDT (Liu et al., 2023f) 2023.02 - SciBERT, ProtBERT SwissProtCLAP $\times$
ProtST-ProtBert (Xu et al., 2023b) 2023.07 420M ProtBert, PubMedBERT ProtDescribe
ProtST-ESM-1b (Xu et al., 2023b) 2023.07 650M ESM-1b, PubMedBERT ProtDescribe
ProtST-ESM-2 (Xu et al., 2023b) 2023.07 650M ESM-2, PubMedBERT ProtDescribe
@@
InstructProtein (Wang et al., 2023c) 2023.10 1.3B OPT (Zhang et al., 2022a) InstructProtein $\times$
@@
DeepTarget (Chen et al., 2023b) 2023.03 - Transformer ChEMBL
@@
BioTranslator (Xu et al., 2023a) 2023.02 - PubMedBERT -
ChatDrug (Liu et al., 2023c) 2023.05 - ChatGPT -
@@ 7.1.1. Molecule-Text Models
Text2Mol (Edwards et al., 2021), KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), MolFM (Luo et al., 2023b), GPT-MolBERTa (Balaji et al., 2023) and MoleculeSTM  (Liu et al., 2023a) represent a suite of encoder-based molecule-text multimodal models. Text2Mol (Edwards et al., 2021) is a multimodal embedding approach to establish an aligned semantic space between text and molecules, facilitating cross-modal retrieval. This approach introduces the CHEBI-20 dataset and employs cross-modal attention-based association rules to enhance the performance in downstream tasks. KV-PLM (Zeng et al., 2022) represents a cross-modal learning approach that integrates molecular structural information with SMILES into textual knowledge. Similar to BERT’s pre-training in an unsupervised fashion, this method facilitates the acquisition of meta-knowledge regarding structural fragments within the textual context. MoMu (Su et al., 2022) is a molecular multimodal model pretrained on both molecular graphs with Graph Isomorphism Network (GIN) (Xu et al., 2018) and their semantically related textual data with SciBERT (Beltagy et al., 2019) via contrastive learning, establishing a crucial link between molecular graphs and natural language. MolFM (Luo et al., 2023b) serves as a multimodal molecular foundation model, facilitating joint representation learning from molecular structures, descriptive texts, and knowledge graphs. It proposes a cross-modal attention mechanism that enhances the comprehension of molecular structures by incorporating atoms, neighboring molecule entities, and semantically related texts. GPT-MolBERTa (Balaji et al., 2023) utilizes ChatGPT to generate rich textual descriptions of molecules, which is then used to pre-train the RoBERTa model in a self-supervised manner. MoleculeSTM (Liu et al., 2023a) employs a molecule encoder and a text encoder with a contrastive learning strategy to jointly learn the chemical structures and their textual descriptions.
@@
DrugChat (Liang et al., 2023), MolXPT (Liu et al., 2023e) and MolReGPT (Li et al., 2023d) are decoder-based molecule-text multimodal models. DrugChat (Liang et al., 2023) combines a graph neural network (GNN) and Vicuna (Chiang et al., 2023) (a classical LLM) to enable ChatGPT-like capabilities on drug molecule graphs. MolXPT (Liu et al., 2023e) is a generative pre-trained model that jointly models text and molecules. It aims to enhance the interaction between molecules and text, facilitating tasks such as molecular property prediction and text-guided molecule generation. MolReGPT (Li et al., 2023d) framework utilizes in-context few-shot molecule learning to enable language models to perform molecule-caption translation tasks without fine-tuning. MolT5 (Edwards et al., 2022), Text+Chem T5 (Christofidellis et al., 2023), ChatMol (Zeng et al., 2023b) and GIT-Mol (Liu et al., 2023b) use encoder-decoder architectures. MolT5 (Edwards et al., 2022) is a T5-based self-supervised learning framework that allows for the translation between molecules and natural language. Text+Chem T5 (Christofidellis et al., 2023) leverages multi-task learning to unify molecular and textual representations, which can solve language tasks, chemical tasks, and cross-domain tasks, without the need for task-specific fine-tuning or retraining. ChatMol (Zeng et al., 2023b) is initialized with T5 and conducts multi-task pre-training, facilitating conversational molecular design through natural language interaction. Based on Q-Former design from BLIP2 (Li et al., 2023c), GIT-Mol (Liu et al., 2023b) proposes a novel modality mixer designed with a cross-attention mechanism, enabling seamless fusion of multiple molecular modalities, including SMILES, texts, images, and structure graphs.
@@ 7.1.2. Protein-Text Models
The utilization of textual data from literature and databases has become increasingly important in protein research, leading to the emergence of protein-text multimodal models. ProTranslator (Xu and Wang, 2022) innovatively approaches protein function prediction by redefining it as a machine translation problem. By utilizing textual descriptions, ProTranslator can annotate proteins to functions without the need for pre-known associated proteins. ProteinDT (Liu et al., 2023f) adopts contrastive learning to align the representation space of text and protein sequences. The ProteinDT model addresses the problem of aligning text and protein sequences to enable text-guided protein design and editing tasks. ProtST series (Xu et al., 2023b), including ProtST-ESM-1B, ProtST-ESM-2, and ProtST-ProtBert, are models enhancing protein sequence pre-training by integrating biomedical texts. Leveraging tasks like unimodal mask prediction and multimodal representation alignment, these models demonstrate superior performance in representation learning benchmarks. Prot2Text (Abdine et al., 2023) predicts a protein’s function in free text format by integrating GNNs and LLMs in an encoder-decoder framework. InstructProtein (Wang et al., 2023c) stands out as a decode-only LLM bridging the gap between human and protein languages through bidirectional generative capabilities.
@@ 7.1.3. Protein-Molecule Models
Protein-molecule models focus on converting protein information into molecular information and exploring potential associations between the two. Protein-molecular models typically employ an encoder-decoder architecture, where the encoder converts a protein sequence into fixed-length vectors, and the decoder generates molecular information from those vectors. ChemBERTaLM (Uludoğan et al., 2022) uses a pre-trained biochemical language model to initialize a target molecule generation model, aiming to address the challenge of limited data on new protein labeling in drug discovery. DeepTarget (Chen et al., 2023b) is an innovative deep generative model that designs new drug molecules directly from protein amino acid sequences, integrating adversarial networks and contrastive learning to minimize reliance on existing structural information. DrugGPT (Li et al., 2023a) is built upon the GPT framework for ligand design, enabling precise capture of both structural information and chemical principles governing drug molecules. DrugCLIP (Gao et al., 2023b) is a novel approach to virtual screening for potential drugs that uses contrastive learning to align representations of binding protein pockets and molecules. These methods aim to improve the efficiency and accuracy of drug discovery through AI-assisted methods.
@@ 7.1.4. Comprehensive Models
In this survey, the comprehensive multimodal models include text, proteins, molecules, genomes and other scientific languages. Their potential impact spans diverse domains, particularly offering significant support in accelerating fundamental science research. Galactica (Taylor et al., 2022b), BioTranslator (Xu et al., 2023a), ChatDrug (Liu et al., 2023c), BioMedGPT0-10B (Luo et al., 2023c), DARWIN-MDP  (Xie et al., 2023), BioT5 (Pei et al., 2023) and Mol-Instructions (Fang et al., 2023a) are representative comprehensive multimodal models. Galactica (Taylor et al., 2022b) is a scientific language model based on a decoder-only architecture, trained on a massive corpus of 106 billion tokens from various scientific texts and knowledge bases. Its strength lies in its ability to efficiently process extensive scientific datasets, enhancing its potential utility in specific scientific tasks. BioTranslator (Xu et al., 2023a) leverages PubMedBERT (Gu et al., 2021) for modeling textual descriptions and fine-tunes it using ontologies across various domains, showcasing its versatility in integrating diverse modalities for sophisticated biological data analysis. ChatDrug (Liu et al., 2023c) utilizes conversational language models to perform drug editing tasks. It addresses the problem of modifying small molecules, peptides, and proteins to achieve specific properties, such as solubility or secondary structure, by leveraging the knowledge extraction and summarization capabilities of the language model. BioMedGPT-10B (Luo et al., 2023c) is an open multimodal generative pre-trained transformer, to bridge the modality gap in biomedicine by aligning different biological modalities with natural language. It addresses the challenge of effectively communicating with diverse biological modalities through free text, including molecular structures, protein sequences, pathways, and cell transcriptomics. DARWIN-MDP (Xie et al., 2023) is a variant of the DARWIN series that involves 16 FAIR datasets to generate instructions for material and device prediction tasks, such as classification, regression, and design. BioT5 (Pei et al., 2023) is a pre-training framework that leverages structured and unstructured data sources to capture the underlying relations and properties of bio-entities. It aims to address the challenges in cross-modal generation and prediction tasks in the field of computational biology. Mol-Instructions (Fang et al., 2023a) aims to address the challenge of enhancing LLMs’ performance in tasks related to molecules, proteins, and biomolecular text by providing task-specific instructions.

### v2

@@ 7. Multi-modal Scientific Large Language Models
MoleculeSTM (Liu et al., 2023b) 2023.12 252M GraphMVP (Liu et al., 2021b), SciBERT Chemformer (Irwin et al., 2022) PubChemSTM
Atomas (Zhang et al., 2024c) 2024.04 60M T5 PubchemSTM-distll
DrugLLM (Liu et al., 2024b) 2024.05 7B LLaMA ZINC, ChEMBL $\times$
3D-MolT5 (Pei et al., 2024b) 2024.06 255M T5 C4, PubMed Central (Canese and Weis, 2013; White, 2020) $\times$
MolecularGPT (Liu et al., 2024a) 2024.06 7B LLaMA2-chat-7B ChEMBL, QM9
@@
ProteinDT (Liu et al., 2023g) 2023.02 - SciBERT, ProtBERT SwissProtCLAP $\times$
ProtST-ProtBert (Xu et al., 2023c) 2023.07 420M ProtBert, PubMedBERT ProtDescribe
ProtST-ESM-1b (Xu et al., 2023c) 2023.07 650M ESM-1b, PubMedBERT ProtDescribe
ProtST-ESM-2 (Xu et al., 2023c) 2023.07 650M ESM-2, PubMedBERT ProtDescribe
@@
InstructProtein (Wang et al., 2023g) 2023.10 1.3B OPT (Zhang et al., 2022b) InstructProtein $\times$
ProLLaMA (Lv et al., 2024) 2024.02 - LLaMA2 UniRef50
ProtLLM (Zhuo et al., 2024) 2024.02 - LLaMA-7B InterPT (Zhuo et al., 2024)
ProtT3 (Liu et al., 2024d) 2024.05 - ESM2, Galactica Swiss-Prot (Bairoch and Apweiler, 2000), ProteinKG25 (Zhang et al., 2022a)
PAAG (Yuan et al., 2024) 2024.04 1.3B SciBERT, ProtBert ProtAnnotation $\times$
Cell&Text Cell2Sentence (Levine et al., 2023) 2023.09 93M/313M GPT-2 CellTypist (Domínguez Conde et al., 2022), cellxgene (Megill et al., 2021)
CELLPLM (Chang and Ye, 2024) 2023.10 80M Transformer GEO (Clough and Barrett, 2016), HCA, HTCA (Regev et al., 2017), SRT (He et al., 2021a)
GenePT (Chen and Zou, 2023) 2023.10 - GPT-3.5 NCBI (Wheeler et al., 2007)
ChatCell (Pei et al., 2024c) 2024.02 - GPT-4 -
GPTCelltype (Hou and Ji, 2024) 2024.03 - GPT-4 -
CellWhisperer (Schaefer et al., [n. d.]) 2024.03 - Geneformer, BioBERT ARCHS4  (Lachmann et al., 2018) $\times$
LangCell (Zhao et al., 2024b) 2024.06 - Geneformer, BERT, PubMedBERT scLibrary (Zhao et al., 2024b)
@@
DeepTarget (Chen et al., 2023c) 2023.03 - Transformer ChEMBL
@@
BioTranslator (Xu et al., 2023b) 2023.02 - PubMedBERT -
ChatDrug (Liu et al., 2023d) 2023.05 - ChatGPT -
@@
BioBridge (Wang et al., 2023e) 2024.01 - KGs PrimeKG
ChemDFM (Zhao et al., 2024a) 2024.01 13B LLaMa-13B Chemical literature, textbooks
ChemLLM (Zhang et al., 2024b) 2024.02 7B InternLM2 ChemData and Multi-Corpus
Evo (Nguyen et al., 2024) 2024.02 7B StripedHyena OpenGenome
BioT5+ (Pei et al., 2024a) 2024.02 252M T5 PubChem, Uniref50, C4
@@ 7.1.1. Molecule-Text Models
Text2Mol (Edwards et al., 2021), KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), MolFM (Luo et al., 2023b), GPT-MolBERTa (Balaji et al., 2023) and MoleculeSTM  (Liu et al., 2023b) represent a suite of encoder-based molecule-text multimodal models. Text2Mol (Edwards et al., 2021) is a multimodal embedding approach to establish an aligned semantic space between text and molecules, facilitating cross-modal retrieval. This approach introduces the CHEBI-20 dataset and employs cross-modal attention-based association rules to enhance the performance in downstream tasks. KV-PLM (Zeng et al., 2022) represents a cross-modal learning approach that integrates molecular structural information with SMILES into textual knowledge. Similar to BERT’s pre-training in an unsupervised fashion, this method facilitates the acquisition of meta-knowledge regarding structural fragments within the textual context. MoMu (Su et al., 2022) is a molecular multimodal model pretrained on both molecular graphs with Graph Isomorphism Network (GIN) (Xu et al., 2018) and their semantically related textual data with SciBERT (Beltagy et al., 2019) via contrastive learning, establishing a crucial link between molecular graphs and natural language. MolFM (Luo et al., 2023b) serves as a multimodal molecular foundation model, facilitating joint representation learning from molecular structures, descriptive texts, and knowledge graphs. It proposes a cross-modal attention mechanism that enhances the comprehension of molecular structures by incorporating atoms, neighboring molecule entities, and semantically related texts. GPT-MolBERTa (Balaji et al., 2023) utilizes ChatGPT to generate rich textual descriptions of molecules, which is then used to pre-train the RoBERTa model in a self-supervised manner. MoleculeSTM (Liu et al., 2023b) employs a molecule encoder and a text encoder with a contrastive learning strategy to jointly learn the chemical structures and their textual descriptions.
@@
DrugChat (Liang et al., 2023), MolXPT (Liu et al., 2023f), MolReGPT (Li et al., 2023e), DrugLLM (Liu et al., 2024b) and MolecularGPT (Liu et al., 2024a) are decoder-based molecule-text multimodal models. DrugChat (Liang et al., 2023) combines a graph neural network (GNN) and Vicuna (Chiang et al., 2023) (a classical LLM) to enable ChatGPT-like capabilities on drug molecule graphs. MolXPT (Liu et al., 2023f) is a generative pre-trained model that jointly models text and molecules. It aims to enhance the interaction between molecules and text, facilitating tasks such as molecular property prediction and text-guided molecule generation. MolReGPT (Li et al., 2023e) framework utilizes in-context few-shot molecule learning to enable language models to perform molecule-caption translation tasks without fine-tuning. DrugLLM (Liu et al., 2024b) is an open large language model specialized for few-shot molecule generation, designed to predict and create new molecules with desired pharmacochemical properties based on limited examples, thereby advancing the efficiency of drug discovery processes. MolecularGPT (Liu et al., 2024a) is an innovative LLM fine-tuned for few-shot molecular property prediction, achieving remarkable performance by leveraging a hybrid instruction set and demonstrating the potential of LLMs as universal few-shot learners in the molecular domain.
MolT5 (Edwards et al., 2022), Text+Chem T5 (Christofidellis et al., 2023), GIMLET (Zhao et al., 2023a), ChatMol (Zeng et al., 2023b), GIT-Mol (Liu et al., 2023c), Atomas (Zhang et al., 2024c) and 3D-MolT5 (Pei et al., 2024b) use encoder-decoder architectures. MolT5 (Edwards et al., 2022) is a T5-based self-supervised learning framework that allows for the translation between molecules and natural language. Text+Chem T5 (Christofidellis et al., 2023) leverages multi-task learning to unify molecular and textual representations, which can solve language tasks, chemical tasks, and cross-domain tasks, without the need for task-specific fine-tuning or retraining. GIMLET (Zhao et al., 2023a) is a novel model specifically designed for zero-shot learning in molecular property prediction. It distinctively integrates graph and text data using a transformer-based approach, with an emphasis on generalized position embedding. ChatMol (Zeng et al., 2023b) is initialized with T5 and conducts multi-task pre-training, facilitating conversational molecular design through natural language interaction. Based on Q-Former design from BLIP2 (Li et al., 2023c), GIT-Mol (Liu et al., 2023c) proposes a novel modality mixer designed with a cross-attention mechanism, enabling seamless fusion of multiple molecular modalities, including SMILES, texts, images, and structure graphs. Atomas (Zhang et al., 2024c) is a multi-modal molecular representation learning framework that employs hierarchical alignment to jointly optimize molecule understanding and generation tasks, achieving state-of-the-art performance without requiring explicit local annotations. 3D-MolT5 (Pei et al., 2024b) is a pioneering framework that integrates 1D molecular sequences and 3D structures into a unified model, enhancing language models’ capabilities for diverse molecular tasks through innovative 3D molecular tokenization and extensive multi-task pre-training.
@@ 7.1.2. Protein-Text Models
Texts from literature and databases have become increasingly important in protein research, leading to the emergence of protein-text multimodal models. ProTranslator (Xu and Wang, 2022) innovatively approaches protein function prediction by redefining it as a machine translation problem. By utilizing textual descriptions, ProTranslator can annotate proteins to functions without the need for pre-known associated proteins. ProteinDT (Liu et al., 2023g) adopts contrastive learning to align the representation space of text and protein sequences. The ProteinDT model addresses the problem of aligning text and protein sequences to enable text-guided protein design and editing tasks. ProtST series (Xu et al., 2023c), including ProtST-ESM-1B, ProtST-ESM-2, and ProtST-ProtBert, are models enhancing protein sequence pre-training by integrating biomedical texts. Leveraging tasks like unimodal mask prediction and multimodal representation alignment, these models demonstrate superior performance in representation learning benchmarks. Prot2Text (Abdine et al., 2023) predicts a protein’s function in free text format by integrating GNNs and LLMs in an encoder-decoder framework. InstructProtein (Wang et al., 2023g) stands out as a decode-only LLM bridging the gap between human and protein languages through bidirectional generative capabilities. ProLLaMA (Lv et al., 2024) leverages a pre-trained general LLM (i.e., LLaMA2) and transforms it for protein language processing through a two-stage training framework. It introduces a novel approach by incorporating Low-Rank Adaptation (LoRA) for efficient training, allowing the model to retain natural language capabilities while learning protein language. ProtLLM (Zhuo et al., 2024) is a versatile crossmodal LLM, pioneers an interleaved protein-text pre-training approach, empowering it with superior performance on both protein-centric and protein-language tasks through a dynamic protein mounting mechanism and protein-as-word language modeling. ProtT3 (Liu et al., 2024d) introduces a novel framework for protein-to-text generation that empowers language models to understand and generate textual descriptions of proteins by integrating a protein language model through a cross-modal projector, enhancing text-based protein understanding tasks. PAAG (Yuan et al., 2024) pioneers a novel multimodal protein design approach by aligning textual annotations with protein sequences, enabling the controlled generation of proteins with specific domains and properties guided by rich, functional descriptions from protein databases.
7.1.3. Gene/Cell-Text Models
Cellular behavior is fundamentally driven by gene expression, which dictates various cellular functions and activities. Gene/Cell-Text models employ cutting-edge natural language processing to decode the intricate language of cellular transcriptomes. By translating gene expression data into textual descriptions, these models effectively narrate cellular activities and interactions. This approach offers new insights into cell behavior, making complex biological processes more comprehensible by describing cellular functions through their underlying gene expression data. Cell2Sentence (Levine et al., 2023) is a transformative approach that teaches language models to process and generate biologically relevant text, enhancing capabilities in cell generation, complex cell type annotation, and data-driven text generation. CellPLM (Chang and Ye, 2024) is a pre-trained transformer model designed to handle single-cell RNA sequencing (scRNA-seq) and spatial transcriptomics (SRT) data. It leverages spatially-resolved transcriptomic data and a Gaussian mixture prior distribution to effectively capture intricate cell-cell relationships, enabling superior performance in tasks like cell type annotation, spatial transcriptomic imputation, and scRNA-seq denoising. GenePT (Chen and Zou, 2023), leveraging GPT-3.5 model, generates comprehensive embeddings for genes and cells by integrating text descriptions from NCBI and gene expression data, enabling advanced genomic analysis. GPTCelltype model (Hou and Ji, 2024) is an R software package that harnesses the predictive capabilities of the GPT-4 language model for high-fidelity cell type annotation in single-cell RNA sequencing analysis, significantly streamlining the process and reducing the need for manual expertise. CellWhisperer (Schaefer et al., [n. d.]) leverages the Geneformer model for transcriptome encoding and BioBERT for text annotations, enabling capabilities such as free-text cell retrieval and zero-shot cell type classification in single-cell RNA sequencing datasets. LangCell (Zhao et al., 2024b) pioneers an integrative pre-training paradigm, harnessing multifaceted ontological annotations to decode the semantic intricacies within single-cell RNA sequencing data, thereby enhancing the precision of cell identity inference across diverse biological spectra.
@@ 7.1.4. Protein-Molecule Models
Protein-molecule models focus on converting protein information into molecular information and exploring potential associations between the two. Protein-molecular models typically employ an encoder-decoder architecture, where the encoder converts a protein sequence into fixed-length vectors, and the decoder generates molecular information from those vectors. ChemBERTaLM (Uludoğan et al., 2022) uses a pre-trained biochemical language model to initialize a target molecule generation model, aiming to address the challenge of limited data on new protein labeling in drug discovery. DeepTarget (Chen et al., 2023c) is an innovative deep generative model that designs new drug molecules directly from protein amino acid sequences, integrating adversarial networks and contrastive learning to minimize reliance on existing structural information. DrugGPT (Li et al., 2023a) is built upon the GPT framework for ligand design, enabling precise capture of both structural information and chemical principles governing drug molecules. DrugCLIP (Gao et al., 2023b) is a novel approach to virtual screening for potential drugs that uses contrastive learning to align representations of binding protein pockets and molecules. These methods aim to improve the efficiency and accuracy of drug discovery through AI-assisted methods.
@@ 7.1.5. Comprehensive Models
In this survey, the comprehensive multimodal models include text, proteins, molecules, genomes and other scientific languages. Their potential impact spans diverse domains, particularly offering significant support in accelerating fundamental science research. Galactica (Taylor et al., 2022b), BioTranslator (Xu et al., 2023b), ChatDrug (Liu et al., 2023d), BioMedGPT0-10B (Luo et al., 2023c), DARWIN-MDP  (Xie et al., 2023), BioT5 (Pei et al., 2023), Mol-Instructions (Fang et al., 2023a), BioBridge (Wang et al., 2023e), ChemDFM (Zhao et al., 2024a), ChemLLM (Zhang et al., 2024b), Evo (Nguyen et al., 2024) and BioT5+ (Pei et al., 2024a) are representative comprehensive multimodal models. Galactica (Taylor et al., 2022b) is a scientific language model based on a decoder-only architecture, trained on a massive corpus of 106 billion tokens from various scientific texts and knowledge bases. Its strength lies in its ability to efficiently process extensive scientific datasets, enhancing its potential utility in specific scientific tasks. BioTranslator (Xu et al., 2023b) leverages PubMedBERT (Gu et al., 2021) for modeling textual descriptions and fine-tunes it using ontologies across various domains, showcasing its versatility in integrating diverse modalities for sophisticated biological data analysis. ChatDrug (Liu et al., 2023d) utilizes conversational language models to perform drug editing tasks. It addresses the problem of modifying small molecules, peptides, and proteins to achieve specific properties, such as solubility or secondary structure, by leveraging the knowledge extraction and summarization capabilities of the language model. BioMedGPT-10B (Luo et al., 2023c) is an open multimodal generative pre-trained transformer, to bridge the modality gap in biomedicine by aligning different biological modalities with natural language. It addresses the challenge of effectively communicating with diverse biological modalities through free text, including molecular structures, protein sequences, pathways, and cell transcriptomics. DARWIN-MDP (Xie et al., 2023) is a variant of the DARWIN series that involves 16 FAIR datasets to generate instructions for material and device prediction tasks, such as classification, regression, and design. BioT5 (Pei et al., 2023) is a pre-training framework that leverages structured and unstructured data sources to capture the underlying relations and properties of bio-entities. It aims to address the challenges in cross-modal generation and prediction tasks in the field of computational biology. Mol-Instructions (Fang et al., 2023a) aims to address the challenge of enhancing LLMs’ performance in tasks related to molecules, proteins, and biomolecular text by providing task-specific instructions. BioBridge (Wang et al., 2023e) stands out as an innovative model within the biomedical sphere, adeptly connecting unimodal Foundation Models using Knowledge Graphs. This model exploits the extensive yet fragmented biomedical datasets to propose a unique method for multimodal integration, circumventing the need for broad-scale retraining. ChemDFM (Zhao et al., 2024a), built on the LLaMa-13B model, is trained on chemical literature to assist with molecular recognition, property prediction, and reaction analysis, and to support research through free-form dialogues. ChemLLM (Zhang et al., 2024b) is a specialized large language model for chemistry, based on the InternLM2 model (Cai et al., 2024c). It utilizes chemical data for instruction tuning and has been evaluated by the proposed ChemBench dataset. On the other hand, ChemDFM (Zhao et al., 2024a), built on the LLaMa-13B model, focuses on training with chemical literature to assist in molecular recognition, property prediction, reaction analysis, and support research through free-form dialogues. PharmGPT (Chen et al., 2024b), built on the LLaMa-13B and LLaMa-70B models, is trained using biopharmaceutical and chemical literature to assist in drug discovery, chemical synthesis assistance, and biological pathway analysis. Evo (Nguyen et al., 2024) utilizes the StripedHyena architecture and is trained on the OpenGenome datasets, which include whole prokaryotic genomes. The model has 7 billion parameters and is capable of modeling and generating DNA sequences at scales ranging from molecular to genome-wide. It is trained on 300 billion nucleotides, allowing it to perform zero-shot function prediction across DNA, RNA, and protein modalities. BioT5+ (Pei et al., 2024a) is an extension of the BioT5 framework, it integrates IUPAC names for improved molecular understanding and includes an extensive corpus of bio-text and molecule data from sources such as bioRxiv and PubChem. The model employs multi-task instruction tuning for generality across tasks and introduces a novel numerical tokenization technique for enhanced processing of numerical data.